### v1

Abstract
This work introduces efficient and accurate spectral solvers for nonlocal equations on bounded domains. These spectral solvers exploit the fact that integration in the nonlocal formulation transforms into multiplication in Fourier space and that nonlocality is decoupled from the grid size. As a result, the computational cost is reduced to $\mathcal{O}(N\log N)$ for an $N$-point discretization grid. Our approach extends the spectral solvers developed in [1] for periodic domains by incorporating the two-dimensional Fourier continuation (2D-FC) algorithm introduced in [7]. We evaluate the performance of the proposed methods on two-dimensional nonlocal Poisson and nonlocal diffusion equations defined on bounded domains. While the regularity of solutions to these equations in bounded settings remains an open problem, we conduct numerical experiments to explore this issue, particularly focusing on studying discontinuities.
Keywords: Spectral methods, Fourier continuation, nonlocal equations, peridynamic models, nonlocal Laplace operators, Fourier multipliers, regularity of solutions.
1 Introduction and overview
In this work, we present a spectral numerical method for solving two-dimensional nonlocal equations on bounded domains. In particular, we solve nonlocal Poisson and nonlocal diffusion equations within bounded domains. We consider nonlocal Laplace operators of the form
[…]
where $\delta>0$ is called the horizon, $\beta$ satisfies $\beta<n+2$, and $c^{\delta,\beta}$ is a scaling constant.
Nonlocal operators of this form have their roots in peridynamics [5, 33, 34, 35] and they have been used in different applied settings, such as fractures [20, 6, 12, 26], image processing [18], tumor growth [25], material failure and damage [22, 13, 21], among others. There have been several mathematical and computational results involving nonlocal equations, see for example [11, 16, 17, 19, 28, 29, 9, 27]. On the mathematical side, the work in [2] introduces the Fourier multipliers for nonlocal Laplace operators, studies the asymptotic behavior of these multipliers, and then applies the asymptotic analysis in the periodic setting to prove regularity results for the nonlocal Poisson equation. The Fourier multipliers $m^{\delta,\beta}$ of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ are defined through Fourier transform by
$\widehat{\mathcal{L}^{\delta,\beta}u}=m^{\delta,\beta}\hat{u}.$ (1)
These multipliers are given by the integral representation (2), and they are used to obtain the eigenvalues of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ in the case of periodic domains by evaluating them on a lattice of points (see Section 4.1 of [2]).
[…]
As pointed out in [14, 15], it is challenging to accurately and efficiently compute the eigenvalues of nonlocal operators. The work in [1] simplifies the computation of these multipliers by representing the $n$-dimensional integral in (2) by a hypergeometric function given by
$m^{\delta,\beta}(\nu)=-\|\nu\|^{2}~{}_{2}F_{3}\left(1,\frac{n+2-\beta}{2};2,\frac{n+2}{2},\frac{n+4-\beta}{2};-\frac{1}{4}\|\nu\|^{2}\delta^{2}\right),$ (3)
for any spacial dimension $n$ and any $\beta<n+2$. This reduces the computation of the multipliers in $\mathbb{R}^{n}$ to computing the 1D smooth function in (3).
The work in [30] and [10] apply these Fourier multipliers to study the regularity of solutions for nonlocal diffusion equations and nonlocal wave equations over the space of periodic distributions, respectively. As emphasized in [30], nonlocal diffusion equation satisfies an instantaneous smoothing effect when the integral kernel is singular with $\beta>n$ and a gradual (over time) smoothing effect when $\beta=n$. However, for integrable kernels ($\beta<n$), the nonlocal diffusion equation is non-smoothing. According to the work in [2], the regularity of the solution $u$ for the nonlocal Poisson equation $\mathcal{L}^{\delta,\beta}u=f$ is the same as the regularity of $f$ when $\beta\leq n$, while when $\beta>n$, the solution is $\beta-n$ more regular than $f$. Specifically, if $f\in H^{s}(T^{n})$ for $s\geq 0$, then the solution $u\in H^{s}(T^{n})$ when $\beta\leq n$ and $u\in H^{s+\beta-n}(T^{n})$ when $\beta>n$.
For computational results, on the other hand, many computational techniques and numerical analysis have been developed for solving nonlocal equations, see for example [9, 11, 15, 3, 8, 14, 1, 36]. Spectral methods are accurate and efficient for solving nonlocal equations because they transform the integral in the nonlocal formulation to multiplication in Fourier space. For example, the application of the operator $\mathcal{L}^{\delta,\beta}$ on a function $u$ results in multiplication in Fourier space, as shown in (1), and this runs at a computational cost of $\mathcal{O}(N\log N)$ for an $N$-point discretization grid. For comparison of the efficiency and accuracy of spectral methods versus standard numerical methods such as finite differences, see [1].
Some spectral methods for nonlocal equations have been developed in [14, 1, 23, 15, 36]. The work in [23] introduces a boundary-adapted spectral method for peridynamic diffusion problems with arbitrary boundary conditions. A spectral method for nonlocal equations on periodic domains has also been developed in [1] with high accuracy. One difficulty that arises in adapting the spectral method developed in [1] to solve nonlocal equations on bounded domains is that the solutions are in general nonperiodic, therefore exhibiting the Gibbs phenomenon as the Fourier coefficients of the solution decay slowly. One way around this is to use Fourier continuation. The idea behind Fourier continuation is to reduce the impact of slowly decaying Fourier coefficients of a smooth but nonperiodic function on a given domain by constructing a smooth periodic extension of the function over a larger domain.
Our proposed solvers build on the Fourier spectral method developed in [1] for periodic domains and incorporate the two-dimensional Fourier continuation (2D-FC) algorithm from [7] to solve two-dimensional nonlocal equations on bounded domains. We also use this approach to perform numerical experiments aimed at exploring the regularity of solutions for nonlocal equations, with particular focus on discontinuities. For integrable kernels, numerical simulations indicate that discontinuities may develop in the solution. In contrast, for singular kernels with $\beta>n+1=3$, the solution appear to remain continuous in the interior of the domain, and no discontinuities are observed.
The paper is organized as follows. In Section 2, we describe the two-dimensional Fourier continuation (2D-FC) algorithm developed in [7], along with minor modifications. Section 3 presents our spectral solvers for two-dimensional nonlocal Poisson and nonlocal diffusion equations on bounded domains. In Section 4, we discuss the numerical results and explore the behavior of solutions in various regimes.
2 Two dimensional Fourier continuation
In this section, we provide a description of the 2D-FC algorithm presented in [7]. The description includes a modification of some of the steps in the algorithm to improve efficiency and accuracy. In particular, we modify the steps in computing the continuation values on the Cartesian grids in the extended region. For consistency, we follow a notation similar to that of [7].
2.1 Overview of 1D blending-to-zero algorithm
In the description given in [4], the blending-to-zero algorithm starts by smoothly extending the vector $f=(f_{1},f_{2},\dots,f_{N})\in\mathbb{R}^{N}$ of values of a complex-valued function $f$ over the set of uniformly spaced nodes $\mathcal{D}=\{x_{1},x_{2},\dots,x_{N}\}$ where for a given step size $h>0$, $x_{j}=x_{1}+(j-1)h.$ Without loss of generality, we assume that $x_{1}=0$ and $x_{N}=1$. The extension relies on some number $d$ of near-boundary function values, $\{f_{1},f_{2},\dots,f_{d}\}$ near $x=0$ and $\{f_{N-d+1},f_{N-d+2},\dots,f_{N}\}$ near $x=1$, and it produces a sequence $\{\tilde{f}_{j}\}_{j=-\infty}^{\infty}$ with the property that
$\tilde{f}_{j}=f_{j}\text{ for }j=1,2,\dots,N.$
The extended vector $\tilde{f}$ is then blended to zero on either end by multiplying with a suitable windowing function. A complete description of this algorithm is provided in [4]. In Figure 1, we demonstrate the 1D blending to zero of the nonperiodic function $f:[0,1]\to\mathbb{R}$ given by $f(x)=\sin(15x)e^{x}$.
Figure 1: A demonstration of the 1D blending to zero on a smooth but nonperiodic function $f(x)=\sin(15x)e^{x}$ defined on [0,1]. The black solid points denote the $d$ near-boundary function values used to produce the extension of $f$ denoted by the blue solid points.
2.2 2D-FC algorithm
Let $f:\overline{\Omega}\to\mathbb{C}$ be given, where the domain $\Omega\subset\mathbb{R}^{2}$ has a smooth boundary $\Gamma=\overline{\Omega}\setminus\Omega$. We assume that values of $f$ are known at the uniform Cartesian grid points within $\overline{\Omega}$ and at the grid points on the boundary. The algorithm first produces 1D blending to zero values for $f$ along directions normal to the boundary $\Gamma$, to produce continuation values that vanish on a certain 2D tangential-normal curvilinear grid around $\Gamma$ as depicted in Figure 2. These continuation values are then interpolated onto Cartesian grid points inside the region $V^{+}$ to produce a 2D blending to zero continuation of $f$. Once all the interpolated values have been obtained at all Cartesian grid points around $\Gamma$, the 2D-FC function is then obtained via a 2D Fast Fourier Transform (FFT) given by
$f^{c}(x,y)=\sum_{k=\frac{N_{x}}{2}+1}^{N_{x}/2}\sum_{l=\frac{N_{y}}{2}+1}^{N_{y}/2}\hat{f}^{c}_{kl}\exp\left(\frac{2\pi ikx}{L_{x}}\right)\exp\left(\frac{2\pi ily}{L_{y}}\right),$ (4)
where $L_{x}$ and $L_{y}$ are the periods in $x$ and $y$ directions respectively. The steps of this algorithm are described in the following subsections.
Figure 2: Geometric constructions of the 2D-FC procedure. The interior strip $V^{-}$ and the exterior strip $V^{+}$ are defined in subsection 2.2.1.
2.2.1 Curvilinear grid along a 2D tangential-normal
The curvilinear grids are introduced within the interior and exterior strips $V^{-}$ and $V^{+}$ respectively, as shown in Figure 2. These strips are given by
$V^{-}$ $=$ $\{q(\theta)-n(\theta)\gamma:0\leq\theta\leq 2\pi\text{ and }0\leq\gamma\leq(d-1)k_{1}\}$
$V^{+}$ $=$ $\{q(\theta)+n(\theta)\gamma:0\leq\theta\leq 2\pi\text{ and }0\leq\gamma\leq Ck_{1}\},$
where $q(\theta)=(x(\theta),y(\theta))$ is the parametrization of $\Gamma$, $n(\theta)=(n_{x}(\theta),n_{y}(\theta))$ is the unit normal to the boundary at $q(\theta)$, and $d,C,k_{1}$ represent the number of matching points near the boundary, the number of continuation points, and the step size along the normals respectively. Consider the following uniform discretization of the interval $[0,2\pi]$
$\Theta_{B}=\{\theta_{p}=pk_{2}:0\leq p<B,~{}k_{2}=2\pi/B\}.$ (5)
Then for each $\theta_{p}\in\Theta_{B}$, the curvilinear grid along the corresponding normal within $V^{-}$ and $V^{+}$ are respectively given as
$V^{-}_{\theta_{p}}$ $=$ $\{r_{p,q}=q(\theta_{p})+n(\theta_{p})(q-d+1)k_{1},~{}0\leq q\leq d-1\}$
$V^{+}_{\theta_{p}}$ $=$ $\{s_{p,q}=q(\theta_{p})+n(\theta_{p})qk_{1}/n_{r},~{}0\leq q\leq C_{r}\},$
where $C_{r}=Cn_{r}$ for some integer refinement factor $n_{r}$ (Remark 1). Let $\mathcal{R}=[a_{0},a_{1}]\times[b_{0},b_{1}]$ be the smallest closed rectangle which contains $\Omega\cup V^{+}$ as shown in Figure 2, and define the following Cartesian grid $\mathcal{H}$ on $\mathcal{R}$ with step size $h$, where the continuation function values will be computed.
=$\mathcal{H}=\{z_{ij}=(x_{i},y_{j}):x_{i}=a_{0}+ih;~{}y_{j}=b_{0}+jh,0\leq i<N_{x}\text{ and }0\leq j<N_{y}\}.$
Remark 1.
The refinement factor $n_{r}$ helps to prevent accuracy loss by making the normal-direction grids at the exterior finer than the grids inherent by the blending-to-zero process.
2.2.2 Computing continuation values on $V^{+}_{\theta_{p}}$
For a given $\theta_{p}\in\Theta_{B}$, a continuation of $f$ on $V^{+}_{\theta_{p}}$ is obtained via the blending-to-zero algorithm described in Section 2.1. We first obtain the approximations of $f$ on the set $V^{-}_{\theta_{p}}$, which are indicated by black solid circles in Figure 3, and then use these approximations to obtain a continuation of $f$ via the 1D blending-to-zero procedure along the corresponding normal direction. The values of $f$ on $V^{-}_{\theta_{p}}$ are obtained by a two-step polynomial interpolation scheme, using polynomials of degree $M-1$. This is briefly described as follows:
Consider the case where $|n_{x}(\theta_{p})|\geq|n_{y}(\theta_{p})|$. We first estimate the values of $f$ at each red point, which is the intersection of the normal line and the vertical Cartesian grid lines as shown in Figure 3, by interpolating vertically the function values at the selected $M$ Cartesian points, shown as open circles, onto the red points. Then we interpolate the values of $f$ at the $M$ red points onto the black solid circles, using a polynomial of degree $M-1$.
The case where $|n_{x}(\theta_{p})|<|n_{y}(\theta_{p})|$ is handled in a similar way by replacing the interpolation along vertical Cartesian grid lines with interpolation along horizontal Cartesian grid lines to obtain the function values at the red points, and then interpolating the function values at the red points onto the black solid points.
Figure 3: Interpolation procedure for evaluating values of $f$ on the set $V_{\theta_{p}}^{-}$ for the case $|n_{x}(\theta_{p})|\geq|n_{y}(\theta_{p})|$. The black solid circles indicate the points that define the set $V^{-}_{\theta_{p}}$. The known values of $f$ at the $M$ open circles are used to interpolate the function values at the red stars (intersection points of the normal and the vertical grid lines). The values of $f$ at the red points are then used to obtain the values of $f$ at the black points.
2.2.3 Computing continuation values on the Cartesian grid
Once all the continuation values have been obtained on $V_{\Theta_{B}}^{+}=\{V_{\theta_{p}}^{+},\theta_{p}\in\Theta_{B}\}$, we then obtain the continuation of $f$ on the Cartesian grid points within $V^{+}$ using a two-step 1D polynomial interpolation scheme. The 2D-FC algorithm is then completed by evaluating corresponding FCs in (4) by means of 2D FFT. The description of the two-step 1D polynomial interpolation scheme is presented below.
For an arbitrary Cartesian point $Q\in\mathcal{H}\cap V^{+}$, we first obtain the nearest boundary parameter value, $\theta_{p}$, around $Q$ using a proximity map. This proximity map can be obtained by associating with each curvilinear discretization point $s_{p,q}$ the nearest Cartesian point using round operator (which rounds $x$ and $y$ coordinates of $Q$ to the nearest integer), resulting in a set $P_{0}\subseteq(\mathcal{H}\cap V^{+})\times V_{\Theta_{B}}^{+}$ of pairs of points. The set $P_{0}$ is then modified by removing multiple associations for a given Cartesian point, and, by adding a next nearest curvilinear neighbor to the Cartesian points that are unassociated. The algorithm then uses the $M$-nearest boundary parameter values
$S_{\theta_{p}}=\{\theta_{p-K_{l}},\theta_{p-K_{l}+1},\dots,\theta_{p},\theta_{p+1},\dots,\theta_{p+K_{r}}\},$
around $\theta_{p}$ (where $K_{l}+K_{r}+1=M$) to obtain the continuation of $f$ at $Q$. Parameter values $\theta_{k}$ with negative values of $k$ are interpreted by periodicity: $\theta_{k}=\theta_{B+k}$.
For each parameter value $\theta_{j},~{}p-K_{l}\leq j\leq p+K_{r}$, we find the signed perpendicular distance from point $Q$ to the normal vector $n(\theta_{j})$, denoted by $\tau_{j}$ in Figure 4(a). Let $b_{j}$ be the perpendicular projection of $Q$ onto the normal vector $n(\theta_{j})$, denoted as the blue points in Figure 4. Define
$\mathcal{T}(\tau_{j})=f^{c}(b_{j}),~{}p-K_{l}\leq j\leq p+K_{r},$
where $f^{c}(b_{j})$ is the continuation of $f$ at $b_{j}$ obtained via a 1D polynomial interpolation on $M$ blending to zero points along $n(\theta_{j})$ (see Figure 4(b)). Clearly, $\mathcal{T}(0)=f^{c}(Q)$. It then follows that a 1D polynomial interpolation procedure on the function $\mathcal{T}(\tau)$, using polynomial of degree $M-1$, can be used to obtain the desired continuation of $f$ at Cartesian point $Q$.
Finally, by applying a 2D FFT to the continuation function values computed above, we obtain the desired Fourier series expression in (4) for the continuation function.
Figure 4: Interpolation procedure for obtaining continuation values of $f$ on $\mathcal{H}\cap V^{+}$. In (a), the blue-star interpolation points are the perpendicular projections of the Cartesian point $Q$ onto the normal vectors, and the $\tau$’s are the signed perpendicular distances from $Q$ to each normal vector. In (b), the interpolation of continuation values from the curvilinear grids to a point $Q$ on the Cartesian grids.
2.2.4 Parameter description and selection
Here, we briefly describe the parameters used in the 2D-FC algorithm. Additionally, we select some specific values for these parameters to get the best result.
  • $d$: number of near-boundary points used to blend $f$ to zero along each normal (subsection 2.1).
  • $C$: number of continuation points in the 1D blending to zero process (subsection 2.1).
  • $k_{1}$: normal step size (subsection 2.2.1)
  • $k_{2}$: boundary step size (subsection 2.2.1)
  • $h$: step size used to construct Cartesian grid $\mathcal{H}$ (subsection 2.2.1)
  • $n_{r}$: refinement factor used to obtain the curvilinear grids along each normal vector (Remark 1)
  • $N=N_{x}\times N_{y}$: number of points in the uniform Cartesian grid (subsection 2.2.1)
  • $B$: number of boundary discretization points (subsection 2.2.1)
  • $M-1$: the degree of the interpolating polynomial (subsections 2.2.2 and 2.2.3)
For a given step size $h$ used to construct the Cartesian grid $\mathcal{H}$, we choose normal step size $k_{1}$ and boundary step size $k_{2}$ to coincide with $h$. The parameters $C=25$ and $n_{r}=6$ were used in the 1D blending to zero procedure described in Subsection 2.1. Finally, we set the interpolating polynomial degree $(M-1)=d$ for all matching point numbers $d$ considered except the interpolation-degree experiments presented in Example 1.
In the following example, we demonstrate the performance of the 2D-FC method using three interpolation degrees: $M-1=d,$ $d+1$, and $d+2$ for a given matching point number $d$.
Example 1.
We consider the FC approximation of the function $f:\Omega\to\mathbb{R}$ given by
$f(x,y)=-(x^{8}+y^{8})\sin(8\pi x)\sin(8\pi y)$ (6)
on the unit disk $\Omega=\{(x,y):x^{2}+y^{2}\leq 1\}$ with three different values of $d$: $d=4$, $5$, and $8$; and three different values of $M$: $M=d+1$, $d+2$, and $d+3$.
Figure 5 displays the relative $L^{2}$ errors as functions of the number of spatial discretization, $N_{1D}$ ($N_{x}=N_{y}\approx N_{1D}+2C$), in one direction within $\Omega$. The errors are computed on a Cartesian grid of step size $h/2$ within $\Omega$. The approximations of $f$ on the finer grid are obtained by means of zero padding [31]. For the case $M=d+1$ (Figure 5(a)), we observe fifth, sixth, and ninth orders of convergence for $d=4$, $d=5$, and $d=8$ respectively, showing convergence of order $d+1$. However, for the case $M=d+2$ (Figure 5(b)), we observe order $d+2$ convergence for $d=5$ and $d=8$, and order $d+1$ convergence for $d=4$. For the case $M=d+3$ (Figure 5(c)), we observe order $d+1$ convergence for $d=4$ and $d=5$, and order $d+2$ convergence for $d=8$. The case $M=d+1$, which gives a similar order of convergence ($d+1$) for the three values of $d$ considered, will be used in all of the numerical experiments presented in this work.
(a) $M=d+1$
(b) $M=d+2$
(c) $M=d+3$
Figure 5: Convergence study for the 2D-FC approximation of the function $f$ in (6) with three choices of matching point numbers $d$: $d=4$, $5$, and $8$; and three choices of $M$: $M=d+1$, $d+2$, and $d+3$. The integer $N_{1D}$ is the number of spatial grid points in one spatial direction within $\Omega$.
In the next example, we demonstrate the 2D-FC expansion of a two-dimensional nonperiodic function defined on a two-dimensional domain contained within the curve given by
$x(\theta)=r(\theta)\cos(\theta);~{}y(\theta)=r(\theta)\sin(\theta),$ (7)
where $r(\theta)=5+\cos(7\theta)/2+\sin(4\theta)/3$ and $0\leq\theta\leq 2\pi.$
Example 2.
We choose $d=4$, $M=5$, $h=0.01$, and employ the 2D-FC method to extend the function
$g(x,y)=6+0.5y-0.2(x+1)^{2}+0.6\sin(4.1\sqrt{x^{2}+y^{2}})\cos(3.8(x-y)),$
over the domain defined in (7). Figure 6 displays both the original function and its extension.
Figure 6: Demonstration of the 2D-FC expansion of a nonperiodic function defined on a 2D domain in the settings of Example 2. (a) the original nonperiodic function $g$. The green curve indicates the boundary of the domain. (b) the 2D-FC expansion of the function $g$.
3 Applications of 2D-FC methods
In this section, we present a 2D-FC based method for solving two-dimensional nonlocal Poisson and nonlocal diffusion equations within a bounded domain $\Omega$ and with Dirichlet conditions on the collar region. The collar region is the extension of $\Omega$ by $\delta$ as shown in Figure 7 and is defined by
$\Omega_{I}=\{{\bf{y}}\in\mathbb{R}^{2}\setminus\Omega:\|{\bf{x}}-{\bf{y}}\|\leq\delta,\text{ for some }{\bf{x}}\in\Omega\}.$
(a) $\beta=1.0$
Figure 7: The collar region
A description of our proposed method is given in the following sections.
3.1 Nonlocal Poisson equation
Consider the following 2D nonlocal Poisson equation with Dirichlet conditions on the collar region
$\begin{cases}\mathcal{L}^{\delta,\beta}u(x,y)=f(x,y),&(x,y)\in\Omega\\ u(x,y)=b(x,y),&(x,y)\in\Omega_{I}\end{cases}.$ (8)
Our approach for solving (8) builds on the spectral method for nonlocal Poisson equation on periodic domains by constructing a periodic extension of the problem using the Fourier continuation algorithm. We then solve the extended problem using the iterative linear algebra solver GMRES [32] to obtain the solution within a specified tolerance. This process is detailed below.
We first transform (8) into the following linear system
$Au=g,\quad A=\begin{pmatrix}L\\ B\end{pmatrix},\quad g=\begin{pmatrix}f\\ b\end{pmatrix},$ (9)
where $L$ and $B$ are respectively the associated matrices of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ and the boundary operator $\mathcal{B}$ given by the Dirichlet conditions on the collar region. To apply GMRES to this system, we do not need to compute $A$ and store it in memory. Instead, it is sufficient to express the forward operator $v\mapsto Av$ as a function. The action of the operator $\mathcal{B}$ is a restriction of the values of $v$ to the collar region. We derive the forward operator function using the properties of $\mathcal{L}^{\delta,\beta}$ as follows:
$u\to u_{\text{ext}}\to\widehat{u}_{\text{ext}}\to\widehat{Lu_{\text{ext}}}=m^{\delta,\beta}\widehat{u}_{\text{ext}}\to Lu_{\text{ext}}\to Lu,$ (10)
where $u_{\text{ext}}$ is the extension of $u$ obtained by the 2D-FC method, $\widehat{u}_{\text{ext}}$ is the Fourier transform of $u_{\text{ext}}$ obtained by means of the fast Fourier transform (FFT), $Lu_{\text{ext}}$ is the inverse Fourier transform of $\widehat{Lu_{\text{ext}}}$, and $Lu$ is the restriction of $Lu_{\text{ext}}$ to the domain $\Omega$. We then obtain the solution of (8) by applying GMRES to the system (9) within a specified tolerance.
3.2 Nonlocal diffusion equation
Consider the following 2D nonlocal diffusion equation with Dirichlet conditions on the collar region and with initial condition
$\begin{cases}u_{t}(x,y,t)=\mathcal{L}^{\delta,\beta}u(x,y,t)+s(x,y,t),&(x,y,t)\in\Omega\times(0,T),\\ u(x,y,0)=u_{0}(x,y),&(x,y)\in\Omega,\\ u(x,y,t)=b(x,y),&(x,y)\in\Omega_{I}.\end{cases}$ (11)
Similar to the nonlocal Poisson solver, we first construct the periodic extension of (11) using the 2D-FC algorithm. We then solve the extended problem using the fourth-order Adams-Bashforth method. The process is detailed below.
First, consider the semi-discrete form of (11), where we replace $u$ by a time dependent vector $\vec{u}$ as follows
$\vec{u}^{\prime}=\mathcal{L}^{\delta,\beta}\vec{u}+\vec{s}:=f(\vec{u}).$ (12)
For a given final time $T>0$ and a time step size $\tau$, we first obtain the vectors $\vec{u}_{1}$, $\vec{u}_{2}$, and $\vec{u}_{3}$, using the fourth-order Runge-Kutta method. Then, for $k=4,5,\dots,K$, where $K=T/\tau$, we obtain $\vec{u}_{k}$ by using the fourth-order Adams-Bashforth method given by
$\vec{u}_{k+4}=\vec{u}_{k+3}+\tau\left(\frac{55}{24}f(\vec{u}_{k+3})-\frac{59}{24}f(\vec{u}_{k+2})+\frac{37}{24}f(\vec{u}_{k+1})-\frac{9}{24}f(\vec{u}_{k})\right).$
To obtain $f(\vec{u}_{k})$, $k=0,1,\dots,K$, we apply the 2D-FC method to extend $\vec{u}$ in (12), as follows
$\vec{u}_{k}\to(\vec{u}_{k})_{\text{ext}}\to\widehat{(\vec{u}_{k})}_{\text{ext}}\to\widehat{(\mathcal{L}^{\delta,\beta}\vec{u}_{k})}_{\text{ext}}=m^{\delta,\beta}\widehat{(\vec{u}_{k})}_{\text{ext}}\to(\mathcal{L}^{\delta,\beta}\vec{u}_{k})_{\text{ext}}\to\mathcal{L}^{\delta,\beta}\vec{u}_{k}+\vec{s}_{k}=f(\vec{u}_{k}),$
where $(\vec{u}_{k})_{\text{ext}}$ is the extension of $\vec{u}_{k}$ obtained by the 2D-FC method, $\widehat{(\vec{u}_{k})}_{\text{ext}}$ is the Fourier transform of $(\vec{u}_{k})_{\text{ext}}$ obtained by means of FFT, $(\mathcal{L}^{\delta,\beta}\vec{u}_{k})_{\text{ext}}$ is the inverse Fourier transform of $\widehat{(\mathcal{L}^{\delta,\beta}\vec{u}_{k})}_{\text{ext}}$, and $\mathcal{L}^{\delta,\beta}\vec{u}_{k}$ is the restriction of $(\mathcal{L}^{\delta,\beta}\vec{u}_{k})_{\text{ext}}$ to the domain $\Omega$.
4 Numerical results
This section demonstrates the accuracy and performance of our proposed method for solving nonlocal equations on bounded domains. In particular, we present the convergence study for the spectral solvers described in subsections 3.1 and 3.2 for the 2D nonlocal Poisson and nonlocal diffusion equations, respectively. Additionally, we study the discontinuities in the solutions of these equations. The 2D-FC parameters are as chosen in subsection 2.2.4.
4.1 Accuracy and convergence test for nonlocal Poisson equation
Here, we construct a manufactured solution to study the accuracy and convergence of the nonlocal Poisson equation (8). To this end, let
$u(x,y)=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$
be the solution of (8), where $r_{1}=10.6418$ and $r_{2}=12.6418$. Clearly, $u(x,y)$ is a highly oscillatory nonperiodic function. To obtain the right hand side function, $f(x,y)$, we apply the operator $\mathcal{L}^{\delta,\beta}$ on $u$, using the property of $\mathcal{L}^{\delta,\beta}$ given in (1), with fixed $\delta=0.4$ and three different values of $\beta:$ $1.2$, $2.0$, and $2.5$. These values of $\beta$ represents $\beta<n$, $\beta=n$, and $\beta>n$. A direct computation of the multipliers $m^{\delta,\beta}$ is provided in Table 1.
   $\delta$ $\overset{\beta=1.2}{m^{\delta,\beta}}$ $\overset{\beta=2.0}{m^{\delta,\beta}}$ $\overset{\beta=2.5}{m^{\delta,\beta}}$
$0.4$ $-82.87098585883194$ $-180.5053934013443$ $-387.0397711705603$
Table 1: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=10.6418$ and $r_{2}=12.6418$, with a fixed $\delta=0.4$ and three different values of $\beta$.
We compute these multipliers using the implementation of the ${}_{2}F_{3}$ hypergeometric function provided by mpmath [24], as detailed in [2]. We then use these manufactured solutions to compute the errors as shown in the following example.
Example 3 (Convergence test for the nonlocal Poisson equation).
[…]
$f(x,y)=\mathcal{L}^{\delta,\beta}b(x,y)=m^{\delta,\beta}b(x,y),$
where $r_{1}=10.6418$, $r_{2}=12.6418$, and the multipliers $m^{\delta,\beta}$ are as given in Table 1.
A convergence study for this test case is shown in Figure 8 for $d=4$ and $d=5$, with fixed $M=d+1$. We use the manufactured solution described above to compute the $L^{2}$ errors reported. We observe seventh and eighth orders of convergence for $d=4$ and $d=5$, respectively, showing convergence of order $d+3$ for the three cases of $\beta$ considered. Moreover, we observe no significant difference in the errors reported for $d=4$ and $d=5$ for these numbers of spatial discretizations. Thus, for the rest of this work, we will use $d=4$ for all numerical computations of nonlocal Poisson equations.
(a) $d=4$
(b) $d=5$
Figure 8: Convergence of the 2D-FC based solution of the nonlocal Poisson equation (8) in the settings of Example 3 with fixed $\delta=0.4$, fixed $M=d+1$, and two different values of $d$.
4.2 Accuracy and convergence test for nonlocal diffusion equation
Similarly, we construct a manufactured solution to study the accuracy and convergence for the nonlocal diffusion equation (11). Let
$u(x,y,t)=e^{-2\pi^{2}\kappa t}\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$
be the solution of (11), where $\kappa=0.1$, $r_{1}=r_{2}=15.6455$. Clearly, $u(x,y)$ is an oscillatory non-periodic function. To obtain the heat source, $s(x,y,t)$, we apply operator $\mathcal{L}^{\delta,\beta}$ on $u$, using the property of $\mathcal{L}^{\delta,\beta}$ given in (1), with fixed $\delta=0.3$ and three values of $\beta:1.0$, $2.0$, and $2.5$. Thus,
$s(x,y,t)=u_{t}-\mathcal{L}^{\delta,\beta}u=\left(-2\pi^{2}\kappa-m^{\delta,\beta}\right)u.$ (13)
A direct computation of the multipliers $m^{\delta,\beta}$ is provided in Table 2. We compute these multipliers using the implementation of the ${}_{2}F_{3}$ hypergeometric function provided by mpmath [24], as detailed in [2].
   $\delta$ $\overset{\beta=1.0}{m^{\delta,\beta}}$ $\overset{\beta=2.0}{m^{\delta,\beta}}$ $\overset{\beta=2.5}{m^{\delta,\beta}}$
$0.3$ $-130.16228859689554$ $-321.3202730766787$ $-689.8419741563309$
Table 2: The multipliers of the nonlocal operator $\mathcal{L}^{\delta,\beta}$ applied on $u=\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $r_{1}=r_{2}=15.6455$, with a fixed $\delta=0.3$ and three different values of $\beta$.
Example 4 (Convergence test for nonlocal diffusion equation).
[…]
$x(\theta)=\cos(\theta)+0.35\cos(\theta)-0.35;~{}y(\theta)=0.7\sin(\theta)~{}\text{ for }0\leq\theta\leq 2\pi$ (14)
using step sizes $h=0.02/2^{k},~{}k=0,1,2,3,4$ and time step size $\tau=5.0\cdot 10^{-07}$. We fix $\delta=0.3$, and choose three values of $\beta:$ $1.0$, $2.0$, and $2.5$, and we set the boundary function $b(x,y)=e^{-2\pi^{2}\kappa t}\sin(2\pi r_{1}x)\sin(2\pi r_{2}y)$, where $\kappa=0.1$, $r_{1}=r_{2}=15.6455$, and the heat source, $s(x,y,t)$, is as given in (13).
A convergence study for this test case is shown in Table 3 for $d=4$ and $d=5$, with fixed $M=d+1$. We use the manufactured solution described above to compute the $L^{2}$ errors reported. We observe sixth and seventh orders of convergence for $d=4$ and $d=5$, respectively, showing convergence of order $d+2$ for the three cases of $\beta$ considered. Moreover, as observed in the case of nonlocal Poisson equation, we observe no significant difference in the errors reported for $d=4$ and $d=5$. Thus, for the rest of this work, we will use $d=4$ for all numerical computations of nonlocal diffusion equations.
   $d$ $h$ $\overset{\beta=1.0}{\varepsilon_{2}}$ Order $\overset{\beta=2.0}{\varepsilon_{2}}$ Order $\overset{\beta=2.5}{\varepsilon_{2}}$ Order
4 $2.00\cdot 10^{-2}$ $1.13\cdot 10^{-05}$ —– $5.33\cdot 10^{-05}$ —– $2.48\cdot 10^{-04}$ —–
$1.00\cdot 10^{-2}$ $1.51\cdot 10^{-07}$ $6.2$ $9.45\cdot 10^{-07}$ $5.8$ $6.81\cdot 10^{-06}$ $5.2$
$5.00\cdot 10^{-3}$ $2.07\cdot 10^{-09}$ $6.2$ $1.48\cdot 10^{-08}$ $6.0$ $1.45\cdot 10^{-07}$ $5.6$
$2.50\cdot 10^{-3}$ $1.87\cdot 10^{-11}$ $6.8$ $9.21\cdot 10^{-11}$ $7.3$ $1.22\cdot 10^{-09}$ $6.9$
$1.25\cdot 10^{-3}$ $2.43\cdot 10^{-13}$ $6.3$ $1.29\cdot 10^{-12}$ $6.2$ $2.29\cdot 10^{-11}$ $5.7$
5 $2.00\cdot 10^{-2}$ $7.87\cdot 10^{-05}$ —– $4.10\cdot 10^{-04}$ —– $1.93\cdot 10^{-03}$ —–
$1.00\cdot 10^{-2}$ $8.84\cdot 10^{-07}$ $6.5$ $5.86\cdot 10^{-06}$ $6.1$ $4.16\cdot 10^{-05}$ $5.5$
$5.00\cdot 10^{-3}$ $5.53\cdot 10^{-09}$ $7.3$ $3.30\cdot 10^{-08}$ $7.5$ $3.23\cdot 10^{-07}$ $7.0$
$2.50\cdot 10^{-3}$ $2.99\cdot 10^{-11}$ $7.5$ $1.87\cdot 10^{-10}$ $7.5$ $2.49\cdot 10^{-09}$ $7.0$
$1.25\cdot 10^{-3}$ $3.49\cdot 10^{-13}$ $6.4$ $1.27\cdot 10^{-12}$ $7.2$ $2.15\cdot 10^{-11}$ $6.9$
Table 3: Convergence of the 2D-FC based solution of the nonlocal diffusion equation (11) in the settings of Example 4 with fixed $\delta=0.3$, fixed $M=d+1$, two different values of $d$: $4$ and $5$, and three choices of $\beta$: $1.0$, $2.0$, and $2.5$.
4.3 Study of discontinuities for the nonlocal Poisson equation
In this section, we study the behavior of the discontinuities in the numerical solutions of nonlocal Poisson equations on bounded domains. For the case of integrable kernels ($\beta<2$), we emphasize that if $f$ is discontinuous, then the solution of the nonlocal Poisson equation is also discontinuous and will share the same location of jumps with $f$. However, for the case of singular kernels with $\beta>3$, we form a conjecture that even if $f$ is discontinuous, the solution will be continuous.
Example 5.
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14). We solve the nonlocal Poisson equation (8) over $\Omega$ using $h=0.0025$ with $\delta=0.4$, and three different values of $\beta:$ $1.0$, $2.0$, and $3.0$, and we set $f(x,y)=4$ and $b(x,y)=x^{2}+y^{2}$.
In this example, since for any $\delta>0$ and $\beta<4$, $\mathcal{L}^{\delta,\beta}(x^{2}+y^{2})=4$ for all $x\in\Omega$, then the unique solution for this problem is given explicitly by $u(x,y)=x^{2}+y^{2}$, which is obviously a continuous function. The numerical solution obtained using our solver is consistent with the above observation as illustrated in Figure 9. To emphasize, the numerical solution represent
(a) $\beta=1.0$
(b) $\beta=2.0$
(c) $\beta=3.0$
Figure 9: Simulation results in Example 5 with a fixed $\delta=0.4$ and three different values of $\beta$.
Example 6.
Let $\Omega$ be the unit disk. We solve the nonlocal Poisson equation (8) over $\Omega$ with $h=0.0025$. We choose $\delta=0.2$, four values of $\beta:$ $1.5$, $2.0$, $2.2$, $3.1$, and we set $f(x,y)=|\sin(\pi x)\sin(\pi y)|$ and $b(x,y)=-x^{2}$.
Figure 10 presents the numerical solution for the three choices of $\beta$ considered. We observe that the solution is discontinuous at $\partial\Omega$, and the magnitude of the jump decreases as $\beta$ increases. Specifically, we observe that for $\beta>3$, the magnitude of the jump is small. The maximum and the minimum magnitudes of the jumps at $\partial\Omega$ for all the values of $\beta$ considered are provided in Table 4, and they are computed using the following formula:
magnitude of the $\text{magnitude of the jump}=\left|\widetilde{u}(x_{\theta},y_{\theta})-b(x_{\theta},y_{\theta})\right|,$
where $\widetilde{u}$ is the extension of $u$ from the interior of $\Omega$ to the boundary $\partial\Omega$, $x_{\theta}$ and $y_{\theta}$ are the $x$ and $y$ coordinates of the boundary $\partial\Omega$ at $\theta\in\Theta_{B}$, and $b$ is the given boundary function.
$\beta=1.5$ $\beta=2.0$ $\beta=2.2$ $\beta=3.1$
$\min$ magnitude $0.0538468$ $0.0286598$ $0.0187264$ $6.24\cdot 10^{-05}$
$\max$ magnitude $0.0646947$ $0.0454215$ $0.0372901$ $0.0097378$
Table 4: The minimum and the maximum magnitudes of the jumps at $\partial\Omega$ in the simulation results presented in Figure 10. We observe that the magnitudes of the jumps decrease as $\beta$ increases.
(a) $\beta=1.5$
(b) $\beta=2.0$
(c) $\beta=2.2$
(d) $\beta=3.1$
Figure 10: Simulation results in Example 6 with a fixed $\delta=0.2$ and four different values of $\beta$. We observe that the magnitude of the jump at the $\partial\Omega$ decreases as $\beta$ increases (see Table 4).
The following example considers the case where $f(x,y)$ is discontinuous. As in the periodic case, we expect the solution to be discontinuous at the location where $f$ is discontinuous.
Example 7.
Let $\Omega$ be the kite-shaped domain contained within the curve defined in (14). We solve the nonlocal Poisson equation (8) over $\Omega$ with $h=0.0025$. We choose $\delta=0.5$, four different values of $\beta:$ $1.2$, $2.0$, $2.5$, $3.1$, and we set the boundary function $b(x,y)=x^{2}+y^{2}$ and
$f(x,y)=\begin{cases}80,&\text{ if }x^{2}+4y^{2}<0.2\\ 4,&\text{ otherwise}\end{cases}.$
In Figure 11, we present the numerical solutions for four different values of $\beta$: $1.2$, $2.0$, $2.5$, and $3.1$. We observe that as $\beta$ increases, the magnitudes of the jump discontinuity inside the domain $\Omega$ and at its boundary $\partial\Omega$ decrease. Furthermore, these discontinuities occur at the exact location where $f$ is discontinuous. Specifically, the maximum and the minimum magnitudes of the jumps in the interior of the domain along the ellipse for all the values of $\beta$ considered are given in Table 5, and they are computed using the following formula:
magnitude of the $\text{magnitude of the jump}=\left|\widetilde{u}_{\text{int}}(x_{\theta},y_{\theta})-\widetilde{u}_{\text{ext}}(x_{\theta},y_{\theta})\right|,$
where $\widetilde{u}_{\text{int}}$ is the extension of $u$ from the interior of the ellipse to the boundary of the ellipse, $\widetilde{u}_{\text{ext}}$ is the extension of $u$ from the exterior of the ellipse to the boundary of the ellipse, and $x_{\theta}$ and $y_{\theta}$ are the $x$ and $y$ coordinates of the boundary of the ellipse at $\theta\in\Theta_{B}$.
$\beta=1.2$ $\beta=2.0$ $\beta=2.5$ $\beta=3.1$
$\min$ magnitude $1.3589861$ $0.3237825$ $0.0363031$ $1.75\cdot 10^{-05}$
$\max$ magnitude $1.3681906$ $0.4040726$ $0.0856388$ $0.0067363$
Table 5: The minimum and the maximum magnitudes of the jumps in the interior of $\Omega$ in the simulation results presented in Figure 11. We observe that the magnitudes of the jumps decrease as $\beta$ increases.
(a) $\beta=1.2$
(b) $\beta=2.0$
(c) $\beta=2.5$
(d) $\beta=3.1$
Figure 11: Simulation results in Example 7 with a fixed $\delta=0.5$ and four different values of $\beta$.
To make a connection between the numerical results in this work, in particular Example 7, and previous mathematical analysis in this area, we note that regularity results for the nonlocal Poisson equation on periodic domains have been developed in [2]. However, regularity of solutions for the nonlocal equation on bounded domains is still an open problem.
For the periodic case, Theorem 4.1 in [2] implies that if $f\in L^{2}$, then $u\in H^{\beta-n}$. In particular, in the case when when $n=2$ and $43<\beta<$ 4, then $u\in H^{\beta-2}$, which implies that $u$ is continuous using the embedding theorem. Using intuition from the periodic case combined with the numerical observations in Example 7, we expect that this smoothing behavior as $\beta$ increases would be true in the case of bounded domains as stated in the following conjecture.
Conjecture 1.
Let $u$ be the solution of (8) and assume that $f\in L^{2}(\Omega)$. Then, if $43<\beta<$ 4, then $u\in C(\Omega)$.
4.4 Evolution of discontinuities for nonlocal diffusion equation
In this section, we study the propagation of discontinuity in the solution of nonlocal diffusion equation (11).
Example 8.
Let $\Omega$ be the domain contained within the curve given by
$x(\theta)=r(\theta)\cos(\theta);~{}y(\theta)=r(\theta)\sin(\theta),$ (15)
where $r(\theta)=1.1+\cos(7\theta)/20+\sin(4\theta)/30,$ and $0\leq\theta\leq 2\pi$. We solve the nonlocal diffusion equation (11) over $\Omega$ at four different time instances: $0.02$, $0.03$, $0.04$, and $0.05$ with time step size $\tau=2.5\cdot 10^{-7}$ and spatial step size $h=0.005$. We set the initial condition $u_{0}(x,y)=\sin(2\pi x)\cos(2\pi y)$, the boundary function $b(x,y)=-(1+x^{2}+y^{2})$, and the heat source $s(x,y,t)=0.$
In this case, the corresponding classical heat equation is given by
$\begin{cases}u_{t}(x,y,t)=\Delta u(x,y,t),&(x,y,t)\in\Omega\times(0,T),\\ u(x,y,0)=\sin(2\pi x)\cos(2\pi y),&(x,y)\in\Omega,\\ u(x,y)=-(1+x^{2}+y^{2}),&(x,y)\in\partial\Omega.\end{cases}$ (16)
The first row of Figure 12 demonstrates the numerical solution to the classical diffusion problem (16) at different times $t=0.02$, $0.03$, $0.04$, and $0.05$. The second row of Figure 12, demonstrates the numerical solution for the nonlocal diffusion problem in Example 8 with fixed $\delta=0.2$ and $\beta=1.5$ for the same corresponding times $t=0.02$, $0.03$, $0.04$, and $0.05$. For classical diffusion, we observe an instantaneous smoothing effect at $\partial\Omega$, in particular, the solution is continuous at the boundary. In contrast, the solution to the nonlocal problem is discontinuous at $\partial\Omega$, however, the magnitude of the jump decreases as time increases. We further observe that for both local and nonlocal diffusion, oscillations in the initial data smooth out as time increases.
(a) $t=0.02$
(b) $t=0.03$
(c) $t=0.04$
(d) $t=0.05$
Figure 12: Simulation results in Example 8 with fixed $\delta=0.2$ and two different values of $\beta$: 4.0 (first row) and 1.5 (second row) at four different time instances.
Example 9.
We solve the nonlocal diffusion equation (8) over the domain contained within the curve given in (15) with spatial step size $h=0.005$ and time step size $\tau=2.5\cdot 10^{-7}$ at four different time instances: $t=0.001$, $0.005$, $0.01$, and $0.03$. We set $b(x,y)=0$, the heat source $s(x,y,t)=0$, and the initial condition
$u_{0}(x,y)=\begin{cases}2,&\text{ if }x^{2}+y^{2}<0.3,\\ e^{-x^{4}-y^{4}},&\text{ otherwise}.\end{cases}$
In Figure 13, we present the numerical solutions for a fixed $\delta=0.2$ and two different values of $\beta$: $4.0$ and $1.0$ at different final time. For the case $\beta=4.0$, which corresponds to the classical heat equation, we observe an instantaneous smoothing effect at $\partial\Omega$ compared to the nonlocal counterpart where $\beta=1.0$.
(a) $t=0.001$
(b) $t=0.005$
(c) $t=0.01$
(d) $t=0.03$
Figure 13: Simulation results in Example 9 with fixed $\delta=0.2$ and two different values of $\beta$: 4.0 (first row) and 1.0 (second row) at four different time instances.